1a.All household circuits are wired in parallel. A 1200-W toaster, a 340-W blender, and a 60-W lamp are plugged into the same outlet. (The three devices are in parallel when plugged into the same outlet.) Assume that this is the standard household 120-V circuit with a 15-A fuse.

What is the current drawn by the toaster?

1b. What is the current drawn by the blender in the previous problem?

1c. What is the current drawn by the lamp in the previous problem?
Answer #1

Since all three equipments are in parallel,

So, voltage across them will be same and equals to standard voltage(120 V).

Now, Power is given by,

P = V*I

1a.

here, V = voltage across toaster = Vt = 120 V

Power of toaster = Pt = 1200 W

So, current drawn by the toaster = it = Pt/Vt = 1200/120

it= 10 A

1b.

here, V = voltage across blender = Vb = 120 V

Power of blender = Pb = 340 W

So, current drawn by the blender = ib = Pb/Vb = 340/120

ib = 2.83 A

1c.

here, V = voltage across lamp = Vl = 120 V

Power of blender = Pl = 60 W

So, current drawn by the blender = il = Pl/Vl = 60/120

il = 0.50 A

Since sum of all currents is less than '15 A', So, circuit should not fuse.

Please upvote.